OpenStudy (anonymous):

omar is 3 times as old as jason. henry is 5 years older than jason. if their total age is 80 years old,how much older is omar than henry?

OpenStudy (anonymous):

Let's say Jason's age is x, Omar's age is y and Henry's age is z. Then we have: y = 3x (Omar is three times as old as Jason) z = x + 5 (Henry is five years older than Jason) x + y + z = 80 (Their total age is 80) Now replace y and z in the last equation with the two other equations and you get: x + 3x + (x + 5) = 80 Calculate brackets: x + 3x + x + 5 = 80 Move 5 over and add up the x: 5x = 75 x = 75/5 = 15 So Jason is 15. Now we put x = 15 into the other equations and we get: Omar's age: y = 3*15 = 45 Henry's age: z = 15 + 5 = 20 So Omar is 25 years older than Henry.